OpenStudy (anonymous):

Find integral sign 10/(4+x^2) dx using the substitution x=2 tan theta. (Give your answer in terms of x.)

OpenStudy (anonymous):

\[10\int\frac{1}{4+x^2}dx\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not sure why you need this substitution, since the derivative of \[\tan^{-1}(x)=\frac{1}{1+x^2}\] so you are close already, but we can do it

OpenStudy (anonymous):

so put \[x=2\tan(\theta)\] \[dx=2\sec^2(\theta)d\theta\] and so you get \[\int\frac{\sec^2(\theta)}{4+4\tan^2(\theta)}d\theta\]

OpenStudy (anonymous):

denominator is \[4(1+\tan^2(\theta))=4\sec^2(\theta)\]

OpenStudy (anonymous):

so now your entire integral is \[\frac{1}{4}\int\frac{sec^2(\theta)}{\sec^2(\theta)}d\theta\]

OpenStudy (anonymous):

in other words just \[\frac{1}{4}\int1d\theta\]

OpenStudy (anonymous):

anti-derivative is \[\frac{1}{4}\theta\]

OpenStudy (anonymous):

now we have to go back. \[x=2\tan(\theta)\iff \frac{x}{2}=\tan(\theta)\iff \theta=\tan^{-1}(\frac{x}{2})\]

OpenStudy (anonymous):

hmm i seem to be off by a factor of 2

OpenStudy (anonymous):

oh i am an idiot. \[x=2\tan(\theta)\] \[dx=2\sec^2(\theta)d\theta\] that is why. there should be another 2 out front!

OpenStudy (anonymous):

so stick a 2 out front and cancel to get \[\int\frac{1}{4+x^2}dx=\frac{1}{2}\tan^{-1}(\frac{x}{2})\]

OpenStudy (anonymous):

i guess you have to multiply by 10 so "final answer" is \[5\tan^{-1}(\frac{x}{2})+C\]

OpenStudy (anonymous):

sorry it took a while but these stupid notifications are blocking my view

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

yw hope steps are clear and sorry i messed up with the 2